# FitCF: A Framework for Automatic Feature Importance-guided Counterfactual Example Generation

### v1

@@ 5.2.3 Effect of Label Flip Verification
Model Dataset AG News ($PPL=95.72$) SST2 ($PPL=309.53$)
Approach Method SLFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$ SLFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$
GPT2 Polyjuice - 18.60% 121.76 0.50 29.00% 258.32 0.71
Llama3-8B FIZLE - 93.50% 123.67 0.61 95.50% 202.22 0.52
[…]
ZeroCF IG 95.50% 109.09 0.27 99.50% 222.51 0.42
ZeroCF LIME 97.50% 107.72 0.39 97.00% 264.91 0.42
ZeroCF SHAP 98.00% 99.08 0.27 94.00% 204.76 0.46
[…]
FitCF IG 96.00% 87.67 0.23 100.00% 161.88 0.48
FitCF LIME 95.50% 75.15 0.19 100.00% 151.22 0.48
FitCF SHAP 94.00% 260.57 0.21 100.00% 157.36 0.49
Qwen2.5-32B FIZLE - 49.00% 53.07 1.14 86.80% 167.51 0.66
[…]
ZeroCF IG 51.00% 60.45 0.76 91.00% 222.57 0.64
ZeroCF LIME 56.00% 63.75 0.84 90.50% 576.59 0.62
ZeroCF SHAP 55.50% 61.68 0.79 93.00% 191.00 0.60
[…]
FitCF IG 57.50% 57.01 0.68 90.50% 221.64 0.49
FitCF LIME 56.00% 57.45 0.79 89.50% 174.34 0.52
FitCF SHAP 62.00% 57.64 0.78 89.50% 157.09 0.52
Qwen2.5-72B FIZLE - 21.50% 84.09 0.22 92.00% 257.91 0.43
[…]
ZeroCF IG 24.50% 92.47 0.22 92.00% 281.10 0.46
ZeroCF LIME 23.00% 72.73 0.71 85.00% 289.20 0.30
ZeroCF SHAP 25.00% 73.92 0.74 86.50% 319.60 0.22
[…]
FitCF IG 42.00% 63.54 0.33 95.00% 207.55 0.39
FitCF LIME 45.00% 61.54 0.35 96.50% 240.94 0.41
FitCF SHAP 38.96% 67.28 0.34 96.50% 590.94 0.39
Table 1: Automatic evaluation results of counterfactuals generated by FIZLE, ZeroCF, and FitCF with Llama3-8B, Qwen2.5-32B, and Qwen2.5-72B using Soft Label Flip Rate (SLFR), Perplexity (PPL), and Textual Similarity (TS) on AG News and SST2. Bold faced values indicate for each approach, which feature importance method is the best performing according to the respective metric.
@@ 5.3 Correlation Analysis
6 Results
6.1 Automatic Evaluation
Table 1 demonstrates that our proposed approaches, ZeroCF and FitCF, consistently outperform Polyjuice easily, which exhibits relatively low SLFR. For AG News dataset using Qwen2.5-32B, the edit distance is comparatively higher than that of Polyjuice, and the other baseline, FIZLE, also shows a larger edit distance compared to Polyjuice. For SST2 dataset, Qwen2.5-72B tends to generate counterfactuals that are less natural and fluent when leveraging ZeroCF and FitCF. Interestingly, Llama3-8B, the smallest model among all evaluated LLMs, achieves the best overall performance. In contrast, Qwen2.5-72B generally underperforms compared to both Llama3-8B and Qwen2.5-32B, as Qwen2.5-72B has a stronger capability to discern the underlying context, making it less prone to flipping labels (App. D, Table 10).
Importantly, FitCF emerges as the most effective method for generating high-quality counterfactuals, consistently outperforming both baselines and ZeroCF across all evaluated settings, underscoring its robustness and effectiveness. This demonstrates the advantage of combining feature importance with the counterfactual generation process. Under the FitCF framework, Integrated Gradients and LIME illustrate superior performance in generating counterfactuals compared to the other two approaches.
6.2 Ablation Study
6.2.1 Effect of Important Words
@@
Table 3: Automatic evaluation results of counterfactuals generated by FitCF with Qwen2.5-72B using $c$ demonstrations.
@@ 6.2.2 Effect of Number of Demonstrations
As shown in Table 3, we find that the number of demonstrations plays an critical role in the performance of FitCF. For AG News, SLFR declines precipitously when the number of clusters ($c$) is used as the number of demonstrations (§5.2.2), while the edit distance shows a slight improvement. In comparison, for SST2, the degree of SLFR diminishment is less conspicuous.
@@ 7 Conclusion
We first introduced ZeroCF, an approach that leverages important words derived from feature attribution methods for counterfactual example generation in a zero-shot setting. Building on this, we proposed FitCF, a framework that automatically constructs high-quality demonstrations using ZeroCF, eliminating the need for human-annotated ground truth for counterfactual generation. FitCF validates counterfactuals via label flip verification for their suitability as demonstrations in a few-shot setting. Empirically, FitCF outperforms two baselines Polyjuice and FIZLE, and our own ZeroCF. Through ablation studies, we identified the three core components of FitCF - number of demonstrations, important words, and label flip verification - as critical to enhancing counterfactual quality. Moreover, we evaluated the faithfulness of feature attribution scores and found that LIME and Integrated Gradients are the most effective feature importance methods for FitCF, consistently producing the most faithful feature attribution scores. Finally, our analysis revealed a strong correlation between the faithfulness of feature attribution scores and the quality of the generated counterfactuals.
@@
The deployed BERT models perform well on fine-tuned tasks (Table 6). However, the LLMs used are not as effective as classifiers compared to BERT models (Table 10) Shin et al. (2020). The quality of the generated counterfactual examples may be affected by the fact that, given an instance, LLMs perceive the label as flipped, even though the actual label is not flipped.
@@
Future work includes investigating the correlation between additional dimensions of feature attribution scores, such as plausibility, coherence and insightfulness, and the quality of counterfactuals through user studies Domnich et al. (2024). We also plan to explore the potential of language models with architectures beyond encoder-only models as a foundation for feature attributions to be used in ZeroCF and FitCF.
@@ F.2 Inference Time
Table 9 shows inference time for ZeroCF and FitCF using Llama3-8B, Qwen2.5-32B and Qwen2.5-32B on AG News and SST2.
@@ G.1 Classification Performance of LLMs
Table 10 displays the accuracy score and error rate on AG News and SST2 datasets using Llama3-8B, Qwen2.5-32B, and Qwen2.5-72B. Our findings indicate that Qwen2.5-32B demonstrates the best classification performance with the lowest error rate, whereas Llama3-8B has the poorest classification performance. Notably, Qwen2.5-72B is the only LLM that generates predictions outside the predefined labels on SST2.

### v3

@@ 5.3 Correlation Analysis
6 Results
6.1 Automatic Evaluation
Table 1 demonstrates that our proposed approaches, ZeroCF and FitCF, consistently outperform Polyjuice and BAE easily, which exhibit relatively low SLFR. Notably, BAE achieves the lowest edit distance compared to Polyjuice, FIZLE, ZeroCF, and FitCF, as it only replaces masked words based on textual embeddings. For AG News dataset using Qwen2.5-32B, the edit distance is comparatively higher than that of Polyjuice, and the other baseline, FIZLE, also shows a larger edit distance compared to Polyjuice. For SST2 dataset, Qwen2.5-72B tends to generate counterfactuals that are less natural and fluent when leveraging ZeroCF and FitCF. Interestingly, Llama3-8B, the smallest model among all evaluated LLMs, achieves the best overall performance. In contrast, Qwen2.5-72B generally underperforms compared to both Llama3-8B and Qwen2.5-32B, as Qwen2.5-72B has a stronger capability to discern the underlying context, making it less prone to flipping labels (Appendix D).
Importantly, FitCF emerges as the most effective method for generating high-quality counterfactuals, consistently outperforming three baselines and ZeroCF across all evaluated settings, underscoring its robustness and effectiveness. This demonstrates the advantage of combining feature importance with the counterfactual generation process. Under the FitCF framework, Integrated Gradients and LIME illustrate superior performance in generating counterfactuals compared to the other two approaches.
6.2 Ablation Study
6.2.1 Effect of Important Words
@@
Table 3: Automatic evaluation results of counterfactuals generated by FitCF with Qwen2.5-72B using $k$ demonstrations.
@@ 6.2.2 Effect of Number of Demonstrations
As shown in Table 3, we find that the number of demonstrations plays an critical role in the performance of FitCF. For AG News, SLFR declines precipitously when the number of clusters ($k$) is used as the number of demonstrations (§5.2.2), while the edit distance shows a slight improvement. In comparison, for SST2, the degree of SLFR diminishment is less conspicuous.
@@ 7 Conclusion
We first introduced ZeroCF, an approach that leverages important words derived from feature attribution methods for counterfactual example generation in a zero-shot setting. Building on this, we proposed FitCF, a framework that automatically constructs high-quality demonstrations using ZeroCF, eliminating the need for human-annotated ground truth for counterfactual generation. FitCF validates counterfactuals via label flip verification for their suitability as demonstrations in a few-shot setting. Empirically, FitCF outperforms three state-of-the-art baselines Polyjuice, BAE and FIZLE, and our own ZeroCF. Through ablation studies, we identified the three core components of FitCF - number of demonstrations, important words, and label flip verification - as critical to enhancing counterfactual quality. Moreover, we evaluated the faithfulness of feature attribution scores and found that LIME and Integrated Gradients are the most effective feature importance methods for FitCF, consistently producing the most faithful feature attribution scores. Finally, our analysis revealed a strong correlation between the faithfulness of feature attribution scores and the quality of the generated counterfactuals.
Future work includes investigating the correlation between additional dimensions of feature attribution scores, such as plausibility, coherence and insightfulness, and the quality of counterfactuals through user studies Domnich et al. (2025). We also plan to explore the potential of language models with architectures beyond encoder-only models as a foundation for feature attributions to be used in ZeroCF and FitCF. Furthermore, we would like to explore counterfactual example generation beyond typical text classification tasks to include, for instance, (long-form) question answering and open-ended mathematical problem, which are particularly challenging even for LLMs Dehghanighobadi et al. (2025).
@@
For feature attribution score evaluation, we focus only on faithfulness, although ferret also supports the evaluation of plausibility. However, plausibility evaluation requires human annotations, so we have left it out and consider it as future work.
Acknowledgment
We are indebted to the anonymous reviewers of ACL 2025 for their helpful and rigorous feedback. This work has been supported by the German Federal Ministry of Research, Technology and Space as part of the projects VERANDA (16KIS2047) and BIFOLD 24B.
@@ F.2 Inference Time
Table 9 shows inference time for ZeroCF and FitCF using Llama3-8B, Qwen2.5-32B and Qwen2.5-32B across AG News and SST2 datasets.
@@ G.1 Classification Performance of LLMs
Table 10 displays the accuracy score and error rate on AG News and SST2 datasets using Llama3-8B, Qwen2.5-32B, and Qwen2.5-72B. Our findings indicate that Qwen2.5-32B demonstrates the best classification performance with the lowest error rate, whereas Llama3-8B has the poorest classification performance. Notably, Qwen2.5-72B is the only LLM that generates predictions outside the predefined labels on SST2. Moreover, as selected LLMs and BERT perform similarly on the two datasets (Table 6), we can assume that the tendency of SLFR will be consistent, and our conclusion should remain unchanged.